OpenStudy (anonymous):

#2- For the curve given by 4x2+y2=48+2xy show that dy/dx=y-4x/y-x #3- For the curve given by 4x2+y2=48+2xy, find the positive y coordinate given that the xcoordinate is 2. #4- For the curve given by 4x2+y2=48+2xy, show that there is a point P with x-coordinate 2 at which the line tangent to the curve at P is horizontal.

OpenStudy (solomonzelman):

1) take the derivative of \(\large\color{black}{ 4x^2+y^2=48+2xy }\) (don't forget the \(\large\color{black}{ dy/dx }\) for each derivative of y) 2) Rearrange the equation in terms of \(\large\color{black}{ dy/dx }\)

OpenStudy (solomonzelman):

this is for number 2.

OpenStudy (solomonzelman):

can you find the derivative of \(\large\color{black}{ 4x^2+y^2=48+2xy }\) ?

OpenStudy (anonymous):

i only need #4 i figured out the others :) thank u so much but can u please help me with #4

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

so you want to show that the horizontal tangent of \(\large\color{black}{ 4x^2+y^2=48+2xy }\) is at x=2.

OpenStudy (solomonzelman):

firstly, you need the derivative (I disconnected sorry)

OpenStudy (solomonzelman):

can you tell me what will the \(\large\color{black}{ dy/dx }\) be?

OpenStudy (anonymous):

can u show me all the steps? that's how i learn math & i'll understand through all of the steps :)

OpenStudy (solomonzelman):

sure

OpenStudy (solomonzelman):

~ find the derivative ~ solve for the derivative ~ the derivative of the function is it's slope, so set y' (or dy/dx, whatever notation is) =0 ~ if you get x=2 after you set y'=0, then your statement in number 4 is true.

OpenStudy (solomonzelman):

so can you perform the steps with me now?

OpenStudy (solomonzelman):

okay, when I differentiate, I get (using y' as my notation) \(\large\color{black}{ 8x+2y{\tiny~}y'=2y+2x{\tiny~}y' }\) \(\large\color{black}{ 2y{\tiny~}y'-2x{\tiny~}y'=2y-8x }\) \(\large\color{black}{{\tiny~}y'(2y-2x)=2y-8x }\) \(\large\color{black}{{\tiny~}y'=\frac{\LARGE 2y-8x }{\LARGE 2y-2x} }\) \(\large\color{black}{{\tiny~}y'=\frac{\LARGE y-4x }{\LARGE y-x} }\)

OpenStudy (solomonzelman):

yes, so what the derivative is said to be in your question, is correct.

OpenStudy (solomonzelman):

then: \(\large\color{black}{{\tiny~}0=\frac{\LARGE y-4x }{\LARGE y-x} }\)

OpenStudy (solomonzelman):

you need f(2) for this, to plug in the point (2,f(2))

OpenStudy (solomonzelman):

\(\large\color{black}{4x^2+y^2=48+2xy }\) \(\large\color{black}{4(2)^2+y^2=48+2(2)y }\) \(\large\color{black}{16+y^2=48+4y }\) \(\large\color{black}{y^2-4y=32 }\) \(\large\color{black}{y^2-4y+4=32+4 }\) \(\large\color{black}{(y-2)^2=36 }\) \(\large\color{black}{y=2\pm6 }\) \(\large\color{black}{y=8,~-4 }\)

OpenStudy (solomonzelman):

so there are actually 2 horizontal tangents, y=8 and the other y=-4.

OpenStudy (solomonzelman):

\(\large\color{black}{{\tiny~}y'=\frac{\LARGE y-4x }{\LARGE y-x} }\) \(\large\color{black}{{\tiny~}0=\frac{\LARGE (8)-4(2) }{\LARGE (8)-(2)} }\) \(\large\color{black}{{\tiny~}0=\frac{\LARGE 0 }{\LARGE 6} }\)

OpenStudy (solomonzelman):

yes there is a horizontal tangent to the curve where x=2, and it is when y=8.

OpenStudy (solomonzelman):

the graph: https://www.desmos.com/calculator/wwlyruii03

OpenStudy (solomonzelman):

I will goof around in an English section for a bit, btu I am out in 10 minutes or so

OpenStudy (solomonzelman):

yeah bye, I got to go in a minute and a half

OpenStudy (anonymous):

thank u so much!!!!!!!!!!!!!

OpenStudy (anonymous):

u went over all of the problem perfectly! i understand it now :) ur awsum thank u